OpenStudy (anonymous):

(-4,-2) and m= 1/2 HELP graphing?

OpenStudy (anonymous):

If that is the point that you are given then you should apply the function m = 1/2 to that point to continue your graph. If m (the slope) = 1/2, then the function will look like this in proper notation: y = x/2

OpenStudy (anonymous):

So, for every value of x that you input, the y value will be 1/2 of that x value. Here, I'll draw it for you:

OpenStudy (anonymous):